# DiffDA: a Diffusion model for weather-scale Data Assimilation

### v2

@@ 2.5 Selection of diffusion model structure
Our method provides a lot of flexibility on the selection of the neural network structure for $\epsilon_{\theta}(\mathbf{x}^{j}_{i},\hat{\mathbf{x}}_{i},j):\mathbb{R}^{n}\times\mathbb{R}^{n}\times\mathbb{N}^{+}\rightarrow\mathbb{R}^{n}$ as any neural network that matches the function signature will work. However, $n$ can be tens of millions in practice, rendering many neural network architectures infeasible due to resource constraints. A neural network must utilize the spatial information in the state vector $\mathbf{x}$ to learn efficiently.
Instead of creating a new architecture, we employ a proven architecture in ML weather forecast models. It has a similar signature $\mathbb{R}^{c\cdot n}\rightarrow\mathbb{R}^{n},c\in\mathbb{N}^{+}$ as the diffusion model. Due to this similarity, neural networks that perform well in forecasting should also do well in the data assimilation tasks, and it is likely to take fewer training steps using pre-trained weights of the forecast model when training the diffusion model. Moreover, we can easily update the backbone of the diffusion model with the state-of-the-art weather forecast model.
@@ 3.1 Implementation
We demonstrate our method in a real-world scenario containing 6 pressure-level variables and 4 surface variables, with a horizontal resolution of ${}^{\circ}$ and 13 vertical levels. This matches the resolution of the WeatherBench2 dataset [Rasp et al., 2023] used by state-of-the-art ML weather forecast models. We use the GraphCast model as the backbone of the diffusion model because the pretrained model takes in states at two consecutive time steps $\mathbf{x}_{i-1},\mathbf{x}_{i}$ to predict $\mathbf{x}_{i+1}$. It takes much less effort than other forecast models to re-purpose it to $\epsilon_{\theta}(\mathbf{x}^{j}_{i},\hat{\mathbf{x}}_{i},j)$ given that there is one-to-one matching between $\mathbf{x}_{i-1}$ and $\mathbf{x}^{j}_{i}$, as well as $\mathbf{x}_{i+1}$ and $\hat{\mathbf{x}}_{i}$. As is determined by the pre-trained GraphCast model, the input size $n$ is set to $(6\times 13+4)\times 721\times 1440$.
@@ 3.2 Training data
We use the WeatherBench2 dataset as the first part of the training data. The dataset contains values for our target atmospheric variables from 1979 to 2016 with a time interval of 6 hours extracted from the ERA5 reanalysis dataset. Data from 1979 to 2015 is used for training, while the rest is used for validation and testing. The second part of the training data is generated by running 48-hour GraphCast predictions (with 8 time steps) using the data from the first part as initial conditions. Then, the two parts are paired up according to their physical time. Before feeding it to the model, the input data is normalized using the vertical-level-wise means and standard deviations provided by the pre-trained GraphCast model. The output of the model is de-normalized using the same means and standard deviations. Since the predicted state $\hat{\mathbf{x}}$ is close to the ground truth $\mathbf{x}^{*}$, we add a skip connection from $\hat{\mathbf{x}}$ to the de-normalized model output and use the normalized difference $\mathbf{x}^{0*}_{i}-\hat{\mathbf{x}},\hat{\mathbf{x}}$ as the actual input data to the diffusion model. Following GraphCast, the normalization and de-normalization use different sets of means and standard deviations for original data $\hat{\mathbf{x}}$ and delta data $\mathbf{x}^{0*}_{i}-\hat{\mathbf{x}}$.
@@ 3.3 Treatment of conditioning for sparse observations
Acknowledging the multidimensional nature of the state vector and that most meteorological observations are co-located horizontally (longitude and latitude), we opt for a simplified setting in the conditioning of sparse observations. In this scenario, the observational data is $m$ sampled columns of the ground truth state vector with $6\times 13+4$ values in each column: $\mathbf{y}\in\mathbb{R}^{(6\times 13+4)\times m}$. The mask is simplified to a 2D mask $\mathbf{m}_{s}\in\mathbb{R}^{721\times 1440}$ which is broadcast to other dimensions when doing point-wise multiplication with the state vector in Equation (missing) 4. Interpolation is also applied independently over 2D horizontal slices for each variable and level. To interpolate appropriately for values on a sphere, we project the 2D slice into 3D space on a unit sphere and use the unstructured interpolation function griddata from the scipy package [Virtanen et al., 2020].
@@ 3.4 Experiment settings and results
Figure 4: Overview of the experiment settings. Hexagons represent atmosphere states, black arrows represent data assimilation, brown solid arrows represent 6-hour prediction, brown dashed arrows represent 48-hour prediction, hexagons with dashed edges and sparse points represent sparse observations, and wide arrows point out targets and references to compare in each experiment.
We demonstrate the effectiveness of our method by performing real-world inspired experiments with increasing complexity (Figure 4). In the basic scenario, we perform data assimilation on 48-hour forecast and observation data, then directly compare the assimilated data with the ground truth. Furthermore, we evaluate our method with autoregressive data assimilation, where the predicted state is produced by a 6-hour prediction from the assimilated state 6 hours ago. We designed this experiment to test whether the assimilated data will deviate from the ground truth data, which is crucial in real-world applications. We also designed another experiment to compare the 48-hour forecast based on assimilated data and ground truth data and evaluate the effect of the data assimilation method on the forecast skill. For single-step data assimilation and forecast on single-step data assimilation, we run 16 parallel experiments with different inputs and random seeds and present the result in average values. For the autoregressive data assimilation, we run each experiment once due to the limitations of computation resources. In all the experiments, observations are simulated by taking random columns from the ERA5 dataset considered as the ground truth. We vary the number of observed columns $m$ in the experiments to test the convergence property of our data assimilation method.
@@
In this experiment, we perform data assimilation from a 48-hour prediction using GraphCast and columns of observations ranging from 1,000 to 40,000, then calculate the error between the assimilated data and the ERA5 data. Finally, we compare this error against GraphCast forecast errors with 6-hour to 48-hour lead times.
The result is presented in Figure 5 where we pick three representative variables closely related to forecast skills [Ashkboos et al., 2022, Rasp et al., 2020] including geopotential at 500hPa (z500), temperature at 850hPa (t850), and temperature at 2m (t2m) as the target variables. More results can be found in subsection A.2. With only 1,000 observed columns ($<0.1\%$ total columns), the assimilated data achieves lower RMSEs than the input 48-hour forecast for z500 and t850. At the same time, the errors of t2m are even lower than the 6-hour forecast. The errors further decrease as the number of observed columns increases. On the other side of the spectra, with 40,000 observed columns ($<3.9\%$ total columns), the RMSEs of all three variables are lower than the 6-hour forecast error. This indicates our assimilation method can improve over the given predicted state while remaining consistent with the observations.
@@
In an operational data assimilation system, data assimilation is performed in an autoregressive way: assimilated data from previous time steps are used for producing forecast at the next time step, which is in turn used for assimilating data at that time step. Following this setting, we perform an autoregressive data assimilation cycle: it starts from a 48-hour forecast at 0 hour, and repeats the data assimilation - 6-hour forecast cycle autoregressively. We train two separate diffusion models for this task, one dedicated to 48-hour forecast inputs and the other dedicated to 6-hour forecast inputs. In the data assimilation cycle, two strategies of sampling the simulated observation columns are used. One is to sample the columns at the same locations (Figure 6). The other resamples data at different locations in each iteration (Figure 7). The strategies help estimate the performance in real-world cases since the reality is the mix of the two strategies where some observations are measured at the same locations, such as weather stations and geostationary satellites, while locations of other observations change with time. We evaluate the autoregressive data assimilation result similar to the previous experiment.
We observe distinct patterns for fixed and non-fixed sampling strategies in the results. With fixed sampling (Figure 6), the initial five iterations show promising performance, but errors begin to accumulate in subsequent iterations. The assimilated data with 1,000 columns of observations ($<0.1\%$ total columns) quickly diverts from the ERA5 data especially for the z500 variable, but this is not the case with more observations. It implies that the given observations are not sufficient to balance the error accumulation in the forecast model. In contrast, the non-fixed sampling strategy shows a slower error accumulation rate than fixed sampling (Figure 7). However, an issue is identified in both strategies: the assimilated errors (solid lines) consistently exceed the interpolation errors (dashed lines) after several iterations. This is less of a concern in real-world scenarios where lots of the observations, such as satellite imagery and radar soundings, can not be explicitly interpolated into atmospheric variables. Notably, the assimilation errors of t2m indicate a diurnal pattern with a period of 24 hours. Since the interpolation errors remain stable over iterations, this diurnal pattern is likely introduced by the GraphCast forecast model.
@@
Since assimilated data is often used as the input of weather forecasts, it is important to test the forecast error using our assimilated data. In this experiment, we perform data assimilation at 0 hour, perform a 48-hour prediction using GraphCast, and then compare the resulting errors against the forecast errors with varying lead times using ERA5 at 0 hour as inputs.
@@
As the scoreboard plot in Figure 8 shows, results from this experiment reveal that the 48-hour forecast errors using assimilated data inputs gradually converge to the 48-hour forecast errors using ERA5 inputs. More results can be found in subsection A.3. The errors using assimilated data cannot be lower than using ERA5 inputs because the ERA5 data is used to simulate observations in the experiment. However, this setup allows us to compare the forecast errors against ones with longer lead times to determine the extent of lead time lost due to the use of assimilated data as forecast inputs. With 8,000 observed columns, the errors for z500, t850, and t2m are lower than the 72-hour forecast errors. This implies that the lead time lost is less than 24 hours for these variables. With 40,000 observed columns, the error for z500 and t850 is approximately equivalent to the 60-hour forecast error. However, the t2m errors remain larger than the 60-hour forecast errors, potentially due to diurnal effects that cause a decrease in the forecast errors of t2m from the 48-hour lead time level to a 60-hour lead time level.
3.5 Limitations and future works
@@
Currently, our method can consume point measurements at a single time step as observations (subsection 2.1). This covers weather station and upper-air balloon measurements but does not include satellite imagery and radar soundings which can be represented as functions of the atmospheric states. We are planning to support broader types of observations and multiple time steps using loss-function-based conditioning at the inference period [Chung et al., 2022, Rozet and Louppe, 2023a]. In addition, we are going to explore adding conditioning of the satellite imagery directly into the diffusion model similar to the conditioning of the predicted state (subsection 2.3). This requires much more training effort but provides more flexibility as the observation function is learned implicitly. With more input observations, we expect to have more accurate and stable results when performing autoregressive data assimilation.
@@
3.6 Conclusion
In summary, we propose a data assimilation method, DiffDA, capable of assimilating high-resolution atmospheric variables up to a horizontal resolution of ${}^{\circ}$, setting a new record for ML-based data assimilation models. Based on the denoising diffusion model, we adapt the pre-trained GraphCast neural network as the backbone model motivated by its compatible input and output shapes. DiffDA’s flexibility also allows for integrating other forecast models, ensuring easy updates and maintenance. A key feature of DiffDA is the enforcement of conditioning on the predicted state during training and inference periods, and the conditioning of sparse observations exclusively at the inference stage. An additional benefit of this conditioning approach is the automatic generation of a post-processing model, should observations not be supplied at inference time.
The experimental results validated the effectiveness of our method: the assimilated data converges to the observations as the number of observed columns increases. With less than $0.77\%$ of total “real-world” data columns and 48-hour forecast, the error of assimilated data is on par with 24-hour forecast error. When used as an input for forecast models, those assimilated data resulted in a maximum lead time loss of 24 hours compared with using the ERA5 dataset as inputs. This enables running data assimilation and simulation in an autoregressive cycle. It remains a challenge to constrain the errors across the autoregressive iterations. This is likely to be solved by assimilating more real-world observations.
All the data assimilation experiments can run on a single high-end PC with a GPU within 15 - 30 minutes per data assimilation step, while a similar task using traditional methods typically requires large compute clusters. This indicates a significant reduction in computational costs. It also opens up the possibility of assimilating more observational data that is otherwise discarded by traditional methods and producing assimilated data with higher accuracy.
4 Ethical and social impact

### v3

@@ 2.5 Selection of Diffusion Model Structure
Our method provides a lot of flexibility in the selection of the neural network structure for $\epsilon_{\theta}(\mathbf{x}^{j}_{i},\hat{\mathbf{x}}_{i},j):\mathbb{R}^{n}\times\mathbb{R}^{n}\times\mathbb{N}^{+}\rightarrow\mathbb{R}^{n}$ as any neural network that matches the function signature will work. However, $n$ can be tens of millions in practice, rendering many neural network architectures infeasible due to resource constraints. A neural network must utilize the spatial information in the state vector $\mathbf{x}$ to learn efficiently.
Instead of creating a new architecture, we adapt a proven ML weather forecasting model architecture. Such models have a similar signature $\mathbb{R}^{c\cdot n}\rightarrow\mathbb{R}^{n},c\in\mathbb{N}^{+}$ as the diffusion model. Due to this similarity, neural networks that perform well in forecasting should also do well in the data assimilation tasks, and it is likely to take fewer training steps using pretrained weights of the forecast model when training the diffusion model. Moreover, we can easily update the backbone of the diffusion model with the state-of-the-art weather forecasting model.
@@ 3.1 Implementation
We demonstrate our method in a real-world scenario containing 6 pressure-level variables (temperature, geopotential, u-wind, v-wind, vertical velocity, specific humidity) and 4 surface variables (2m temperature, 10m u-wind, 10 v-wind, mean sea level pressure), with a horizontal resolution of 0.25 and 13 vertical levels (50hPa, 100hPa, 150hPa, 200hPa, 250hPa, 300hPa, 400hPa, 500hPa, 600hPa, 700hPa, 850hPa, 925hPahPa, 1000hPa). This matches the resolution of the WeatherBench2 dataset [Rasp et al., 2023] used by state-of-the-art ML weather forecasting models. We use the GNN-based GraphCast model as the backbone of the diffusion model because the pretrained model takes in states at two consecutive time steps $\mathbf{x}_{i-1},\mathbf{x}_{i}$ to predict $\mathbf{x}_{i+1}$. It takes much less effort than other forecast models to re-purpose it to $\epsilon_{\theta}(\mathbf{x}^{j}_{i},\hat{\mathbf{x}}_{i},j)$ given that there is one-to-one matching between $(\mathbf{x}_{i-1},\mathbf{x}_{i},\mathbf{x}_{i+1})$ and $(\hat{\mathbf{x}}_{i},\mathbf{x}^{j}_{i}.\mathbf{x}^{j-1}_{i})$. As is determined by the pre-trained GraphCast model, the input size $n$ is set to $(6\times 13+4)\times 721\times 1440$.
@@ 3.2 Training Data
We use the WeatherBench2 dataset as the first part of the training data. The dataset contains values for our target atmospheric variables from 1979 to 2016 with a time interval of 6 hours extracted from the ERA5 reanalysis dataset. The training process uses data from 1979 to 2015. Validation uses data in 2016 and testing uses data in 2022. The second part of the training data is generated by running 48-hour GraphCast predictions (with 8 time steps) using the data from ERA5 as initial conditions. Then, the two parts are paired up according to their physical time. Before feeding it to the model, the input data is normalized using the vertical-level-wise means and standard deviations provided by the pre-trained GraphCast model. The output of the model is de-normalized using the same means and standard deviations. Since the predicted state $\hat{\mathbf{x}}$ is close to the ground truth $\mathbf{x}^{*}$, we add a skip connection from $\hat{\mathbf{x}}$ to the de-normalized model output. For the autoregressive data assimilation experiment, we train an additional data assimilation model dedicated for 6-hour forecasts.
@@ 3.3 Treatment of Conditioning for Sparse Observations
Acknowledging the multidimensional nature of the state vector and that most meteorological observations are co-located horizontally (longitude and latitude), we opt for a simplified setting in the conditioning of sparse observations. In this scenario, the observational data is $m$ sampled columns of the ground truth state vector with $6\times 13+4$ values in each column: $\mathbf{y}\in\mathbb{R}^{(6\times 13+4)\times m}$. We set Gaussian kernel standard deviation $\sigma_{G}$ to 2.5 according to the ablation study and kernel diameter $d$ to $14\sigma_{G}+$ 1 (using number of grid cells as the unit). As kernels are applied to observations at different latitudes, $\sigma_{G}$ is scaled with cosine of latitude according to the sphere geometry. The mask is simplified to a 2D mask $\mathbf{m}_{s}\in\mathbb{R}^{721\times 1440}$ which is broadcast to other dimensions when doing point-wise multiplication with the state vector in Equation (missing) 3. Interpolation is also applied independently over 2D horizontal slices for each variable and level. We use the universal kriging interpolation method provided by the pyinterp package to properly interpolate unstructured data points on the sphere. The data points are subtracted by climatology values from WeatherBench2 to conform to the prerequisites of kriging interpolation.
@@ 3.4 Experiment Settings and Results
Figure 4: Overview of the experiment settings. Single-step data assimilation takes in observations and a 48h forecast, and outputs assimilated data at 0h. Autoregressive data assimilation combines a data assimilation model and a 6h prediction model to produce assimilated data every 6h autoregressively. It is also of interest to perform 48h forecasts on single-step assimilated data. Hexagons represent atmosphere states, black arrows represent data assimilation, brown solid arrows represent 6-hour prediction, brown dashed arrows represent 48-hour prediction, hexagons with dashed edges and sparse points represent sparse observations, and wide arrows point out targets and references to compare in each experiment.
We demonstrate the effectiveness of our method by performing real-world inspired experiments of increasing sophistication (Figure 4). First, in the single step experiment, we perform data assimilation on 48-hour forecast and observation data, then directly compare the assimilated data with the ground truth. Second, we evaluate our method with autoregressive data assimilation, where a 6h prediction and data assimilation cycle is run repeatedly. We design this experiment to test whether the assimilated data will deviate from the ground truth data after several autoregressive iterations, which is crucial in real-world applications. Lastly, we designed an experiment to compare the 48-hour forecast based on assimilated data and ERA5 data and evaluate the effect of the data assimilation method on the forecast skill. For single-step data assimilation and forecast on single-step data assimilation, we run 16 parallel experiments with inputs from different time steps and different random seeds. For each experiment, we calculate the assimilation errors against the ERA5 dataset in root mean square errors (RMSEs) and then present the averaged RMSEs in the results. For the autoregressive data assimilation, we run each experiment once due to the limitations of computation resources. In all the experiments, observations are simulated by taking random columns from the ERA5 dataset considered as the ground truth. We vary the number of observed columns $m$ in the experiments to test the convergence property of our data assimilation method.
@@
In this experiment, we perform data assimilation from a 48-hour prediction using GraphCast and columns of observations ranging from 1,000 to 40,000, then calculate the RMSE errors between the assimilated data and the ERA5 data. Finally, we compare the errors against GraphCast forecast errors with 6-hour to 48-hour lead times.
The result is presented in Figure 5 where we pick three representative variables closely related to forecast skills [Ashkboos et al., 2022, Rasp et al., 2020] including geopotential at 500hPa (z500), temperature at 850hPa (t850), and temperature at 2m (t2m) as the target variables. More results can be found in subsection A.2. Generally, assimilation errors decrease as number of observed columns increases. With 4,000 observed columns ($<0.4\%$ total data), the assimilated data achieves lower RMSEs than the input 48-hour forecast for all three variables. The errors further decrease as the number of observed columns increases. On the other side of the spectra, with 40,000 observed columns ($<3.9\%$ total columns), the RMSEs of z500 and t2m are lower than the 6-hour forecast error.
As a case study, we present the assimilated t850 with 8000 observed columns in Figure 6. The assimilated data have lower errors than the two inputs of data assimilation (48h forecast and interpolated observations). It is also better than simply mixing forecast with interpolated observations. In addition, the assimilated data introduce fine grained details which helps in recreating high-frequency spectral component lost during autoregressive forecast.
Figure 6: Plots of assimilated temperature at 850hPa and errors against ERA5 compared with 48h GraphCast forecast, interpolated observations, and mixture of 48h forecast and interpolated observation according to the soft mask. The assimilated data is better in accuracy than both inputs. It also introduces more fine grained details than 48h forecast.
@@
In an operational system, data assimilation is performed in an autoregressive way: assimilated data from previous time steps are used for producing forecast at the next time step, which is in turn used for assimilating data at that time step. Following this setting, we perform an autoregressive data assimilation cycle: it starts from a 48-hour forecast at 0 hour, performs data assimilation at 0 hour, then performs a 6-hour forecast at 6 hour, and repeats the data assimilation – 6-hour forecast cycle autoregressively. We train two separate diffusion models for this task, one dedicated to 48-hour forecast inputs and the other dedicated to 6-hour forecast inputs. In the data assimilation cycle, two strategies of sampling the simulated observation columns are used. One is to sample the columns at the same locations (Figure 7). The other resamples data at different locations in each iteration (Figure 8). The strategies help estimate the performance in real-world cases since the reality is the mix of the two strategies where some observations are measured at the same locations, such as weather stations and geostationary satellites, while locations of other observations change with time. We evaluate the autoregressive data assimilation result similar to the previous experiment. We compare the assimilation errors against errors of interpolated observations which gives a baseline for assimilated data.
We observe distinct patterns for fixed and non-fixed sampling strategies in the results. With fixed sampling (Figure 7), the errors accumulates as the autoregressive data assimilation proceeds especially for fewer observations cases. The assimilation errors of z500 stabilize after several iterations and remain lower than the interpolation error. Meanwhile, assimilated t850 and t2m start from low errors and their errors grow gradually until exceeding the interpolation errors. In particular, our model struggles to assimilate t2m, where assimilated t2m deteriorate quickly and reach to higher error levels than interpolated observations. In contrast, the non-fixed sampling strategy shows a slower error accumulation rate than fixed sampling (Figure 8). The assimilation errors of z500 and t850 remain lower than interpolation errors of observations throughout the autoregressive assimilation process. However, it is not the case for t2m where assimilation errors exceed interpolation errors after a few iterations. This is likely due to the lack of four dimensional assimilation which can improve accuracy and stabilize the temporal roll out. We are working on four dimensional assimilation extension to our method and eventually target for assimilating real-world observations.
@@
Since assimilated data are often used as the input of weather forecasts, it is important to test the forecast errors using our assimilated data. In this experiment, we perform data assimilation at 0 hour, perform a 48-hour prediction using GraphCast, and then compare the resulting errors against the forecast errors with varying lead times using ERA5 as inputs.
@@
As the scoreboard plot in Figure 9 shows, results from this experiment reveal that the 48-hour forecast errors using assimilated data inputs gradually converge to the 48-hour forecast errors using ERA5 inputs. More results can be found in subsection A.3. The errors using assimilated data cannot be lower than using ERA5 inputs because the ERA5 data are used to simulate observations in the experiment. However, this setup allows us to compare the forecast errors against ones with longer lead times to determine the extent of lead time lost due to the use of assimilated data as forecast inputs. With 10,000 observed columns, the errors for z500, t850, and t2m are lower than the 72-hour forecast errors. This implies that the lead time lost is less than 24 hours for these variables. With 40,000 observed columns, the errors are lower than the 60-hour forecast errors.
4 Ablation Study
We perform an ablation study on the Gaussian kernel standard deviation $\sigma_{G}$ to evaluate the effectiveness of the soft mask technique for incorporating sparse observations. A larger $\sigma_{G}$ corresponds to a larger effective range of the observation values. The soft mask degenerates into a hard mask when $\sigma_{G}=0$. We replicate the experiment settings from ‘single-step data assimilation’ and ‘forecast on single-step data assimilation’, varying $\sigma_{G}$ from $0.5$ to $3.0$.
The ablation results for ‘forecast on single-step data assimilation’ are shown in Table 1. Results for ‘single-step data assimilation’ can be found in Appendix Table 2. We observe an increase in accuracy when $\sigma_{G}$ grows from 0.5, demonstrating the advantage of the soft mask technique. The advantage likely comes from the growing region of the masked areas whereas the small masked areas in the hard mask can be smeared out by down-sampling layers in the diffusion model. The errors saturate and increase again for $\sigma_{G}$ between 2.0 and 3.0 depending on the variable. In particular, t2m favors larger $\sigma_{G}$ than z500 and t850. Note that including too many interpolated values will also include their errors. In our main experiments, we chose $\sigma_{G}=2.5$.
Table 1: Root mean square errors (RMSEs) of geopotential at 500hPa, temperature at 850hPa, and temperature at 2m from the 48-hour forecast using assimilated data as inputs varying $\sigma_{G}$ and number of emulated observed columns.
$\sigma_{G}$
#Col 0.5 1.0 1.5 2.0 2.5 3.0
z500 1000 224 216 209 200 209 212
2000 214 207 191 189 200 200
4000 211 194 177 180 187 181
8000 189 173 162 157 166 151
10000 180 164 148 147 146 142
20000 166 143 132 120 119 130
40000 139 120 115 111 112 105
t850 1000 1.53 1.49 1.46 1.42 1.47 1.46
2000 1.47 1.44 1.40 1.38 1.38 1.43
4000 1.47 1.39 1.32 1.31 1.34 1.32
8000 1.37 1.27 1.22 1.21 1.22 1.21
10000 1.33 1.26 1.18 1.16 1.19 1.18
20000 1.26 1.14 1.11 1.09 1.10 1.12
40000 1.14 1.07 1.07 1.06 1.06 1.04
t2m 1000 1.67 1.64 1.61 1.56 1.60 1.55
2000 1.63 1.59 1.55 1.56 1.49 1.55
4000 1.63 1.54 1.48 1.49 1.46 1.47
8000 1.54 1.46 1.42 1.41 1.40 1.41
10000 1.50 1.43 1.42 1.37 1.39 1.40
20000 1.43 1.36 1.35 1.33 1.32 1.33
40000 1.36 1.31 1.31 1.30 1.30 1.29
5 Related Works
Fengwu-4DVar [Xiao et al., 2024] achieves 4-dimensional data assimilation with 1.4 resolution using the traditional 4D variational method. It utilizes the differentiable ML weather forecasting model Fengwu. Score-based data assimilation [Rozet and Louppe, 2023a] utilizes a diffusion model for data assimilation. It uses an additional loss function to enforce the observation values. It has been applied to assimilate 2D geospheric flows  [Rozet and Louppe, 2023b].
MetNet3 [Andrychowicz et al., 2023] can make regional weather forecasts from gridded weather data and sparse observations. It combines data assimilation with prediction into one end-to-end process which requires training with sparse observation data. MetNet3 is susceptible to the generalization problem when training with observation data, thus it is hard to scale to a global high resolution scenario. GenCast [Price et al., 2023] is a diffusion model dedicated to probabilistic global weather forecast. It contains GraphCast-like encoder and decoder, and a transformer based processor. SEEDS [Li et al., 2023] is a diffusion model for generative emulation of weather forecast ensembles. It is capable of post-processing weather forecasts into ensembles which better represent the underlying probabilistic distribution of future weather states.
6 Limitations and Future Works
@@
Currently, our method can consume point measurements at a single time step as observations (subsection 2.1). This covers weather station measurements, upper-air balloon measurements, aircraft measurements, and satellite retrievals but does not include satellite imagery and radar soundings which can be represented as functions of the atmospheric states. We are planning to support broader types of observations and multiple time steps using loss-function-based conditioning at the inference period [Chung et al., 2022, Rozet and Louppe, 2023a]. In addition, we are going to explore adding conditioning of the satellite imagery directly into the diffusion model similar to the conditioning of the predicted state (subsection 2.3). This requires much more training effort but provides more flexibility as the observation function is learned implicitly. With more input observations, we expect to have more accurate and stable results when performing autoregressive data assimilation.
@@
7 Conclusion
In summary, we propose an easy-to-use data assimilation method, DiffDA, capable of assimilating high-resolution atmospheric variables up to a horizontal resolution of 0.25, setting a new record for ML-based data assimilation models. Based on the denoising diffusion model, we adapt the pre-trained GraphCast neural network as the backbone model motivated by its compatible input and output shapes. DiffDA’s flexibility also allows for integrating other forecast models, ensuring easy updates and maintenance. A key feature of DiffDA is the conditioning on sparse observations occurs exclusively during inference which avoids “the curse of dimensionality” of adding sparse observations during training. An additional benefit of this conditioning approach is the automatic creation of a post-processing model, should observations not be supplied at inference time.
The experimental results validated the effectiveness of our method: the assimilated data converge to the observations as the number of observed data points increases. With observations occupying less than $0.96\%$ of total grid points and 48-hour forecast, the errors of assimilated data are on par with 24-hour forecast errors. When used as an input for forecast models, those assimilated data resulted in a maximum lead time loss of 24 hours compared to using ERA5 as inputs. This enables running data assimilation and simulation in an autoregressive cycle. It remains a challenge to constrain the errors across the autoregressive iterations.
All the data assimilation experiments can run on a single high-end PC with a GPU within 15 - 30 minutes per data assimilation step, while a similar task using traditional methods typically requires large compute clusters. This indicates a significant reduction in computational costs. It also opens up the possibility of assimilating more observational data that is otherwise discarded by traditional methods, producing assimilated data with a higher accuracy.
Acknowledgements
This study received funding from the MAELSTROM project funded by the European High-Performance Computing Joint Undertaking (JU; grant agreement No 955513).
Impact Statement